what does it means to raise a power to a power and how it is related to multiplication. Provide an example to illustrate.


Answer by stanbon(48569)

what does it means to raise a power to a power and how it is related to multiplication. Provide an example to illustrate.
----------
(x^2)^3 = x^2*x^2*x^2 = x^6
----------------
(x^-3)^(-2) = (1/x^-3)^2 = 1/[x^-3*x*-3] = 1/x^-6 = x^6
-----------------
General Rule:
(x^n)^m = x^(nm)
----------------
(x^3)^-2 = x^-6 = 1/x^6
=====================
Cheers,
Stan H.